The differential equation of y = a cos (log x) + b sin (log x), where a and b are the parameters is...

Grade:12

1 Answers

Aditya Gupta
4 years ago
we see that dy/dx= y’= a(-sin(logx))/x + bcos(logx)/x
or xy’= – asin(logx)+bcos(logx)
differentiate again
y’+xy”= – acos(logx)/x + b( – sin(logx))/x
or xy’+x^2y”= – a cos (log x) – b sin (log x)= – y
or x^2*d2y/dx2 + xdy/dx + y=0